OpenStudy (anonymous):

how do i factor x^1/2-3x?

OpenStudy (solomonzelman):

You mean \(\normalsize\color{blue}{ x^{1/2}-3x }\) right ?

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

I do not think that you can factor such an expression though...

OpenStudy (solomonzelman):

You can re-write it in terms of a root, \(\normalsize\color{blue}{ \sqrt{x}-3x }\) , and see why not ?

OpenStudy (anonymous):

can i simplify it then?

OpenStudy (solomonzelman):

No, I really don't think that you can do anything with it that would make it look any simpler than what it is now.

OpenStudy (solomonzelman):

I'll post a bit of instructions, okay ?

OpenStudy (anonymous):

Ok

OpenStudy (jdoe0001):

\(\Large \bf { x^{\frac{1}{2}}-3x\implies x^{\frac{1}{2}}-3x^1\implies x^{\frac{1}{2}}-3x^{\frac{2}{2}} \\ \quad \\ x^{\frac{1}{2}}-3x^{\frac{1}{2}+\frac{1}{2}}\implies x^{\frac{1}{2}}-3x^{\frac{1}{2}}x^{\frac{1}{2}} }\) see any common factors?

OpenStudy (anonymous):

yes

OpenStudy (jdoe0001):

so just take that common factor... and is all you could really factor it about, I'd think

OpenStudy (solomonzelman):

\(\LARGE\color{darkblue }{\bf {Also,~~~for~~~a~~~future~~~reference:} }\) \(\normalsize\color{blue}{\bf {you~~~can~~~find~~~any~~~section~~~by:} }\) \(\normalsize \color{black}{\rm { 1) } }\) clicking "more subjects" (the top left corner, towards the center) \(\LARGE\color{black}{\rm { 2) } }\) typing in the name of the section (name of a section can be "history" Mathematics and other section for each unique area of studying). \(\normalsize\color{blue}{\bf {You~~~can~~~FAN~~~someone~~~by:} }\) \(\normalsize\color{black}{\rm { 1) } }\) clicking on the user's name \(\normalsize\color{black}{\rm { 2) } }\) clicking "fan" on the bottom right corner of user's box. \(\normalsize\color{black}{\bf { (~ } }\)You can also write a testimonial by clicking "add testimonial" and typing something in testimonial box.\(\normalsize\color{black}{\rm { ~) } }\) \(\normalsize\color{blue}{\bf {You~~~can~~~mention~~~someone~~~in~~~your~~~question} }\) (or in any other question) by writing @ in front of their name, just like I would mention you... LOOK ! .... @SolomonZelman (they would get a notification that a user mentioned them in a reply) \(\normalsize\color{blue}{\bf {You~~~can~~~ message~~~someone~~~by:} }\) \(\normalsize\color{black}{\rm { 1) } }\) clicking on the someone's username \(\normalsize\color{black}{\rm { 2) } }\) clicking "Send Message" \(\normalsize\color{black}{\rm { 3) } }\) type your message press submit, and wait till the messages submits. \(\normalsize\color{blue}{\bf {You~~~can~~~ see~~~if~~~a~~~user~~~is~~~offline~~~by} }\) Clicking on the user's username, and below his picture you will see (in light blue) “offline”, “online” or you might see that he is viewing or replying to a specific question. You might as well see that he is just looking around in a specific section. If you have any questions you can message me or anyone else on this website. We will be happy to respond !

OpenStudy (solomonzelman):

I meantion myself instead,... but I see you figured out some things already.

OpenStudy (solomonzelman):

Excuse the long reply.

OpenStudy (anonymous):

Its ok thanks so much for the help @SolomonZelman

OpenStudy (solomonzelman):

Anytime !